OpenStudy (anonymous):

implicit differentiation 2x^2+7 x y-y^2=8

OpenStudy (abdullah1995):

take the derivative of 2x^2 using power rule and let me know what you get

OpenStudy (anonymous):

4x

OpenStudy (abdullah1995):

very well. lets move to the next term. 7xy

OpenStudy (anonymous):

7x dy/dx + y?

OpenStudy (abdullah1995):

ok so 7 here is a constant. so you can for now just leave it out for now. focus on xy. the derivative of xy is (1)(y)+(x)(1)(y') ; y' can also be written as dy/dx

OpenStudy (abdullah1995):

then you multiply 7 with the derivative so it becomes 7(y+x dy/dx) or 7(y+xy')

OpenStudy (anonymous):

so then can I also say 7x dy/dx + 7y ?

OpenStudy (abdullah1995):

yes you can

OpenStudy (anonymous):

so far I had -4x-7y/(7x-2y) when I did it myself but answer is showwing wrong.

OpenStudy (anonymous):

apparently my signs are all wrong, not sure where I went wrong

OpenStudy (abdullah1995):

lets move on. take derivative of -y^2

OpenStudy (anonymous):

ok... uh -2y dy/dx ?

OpenStudy (abdullah1995):

yup that's right

OpenStudy (abdullah1995):

and finally derivative of 8 = 0

OpenStudy (anonymous):

so we have 4x+7x dydx + 7y - 2y dydx=0

OpenStudy (abdullah1995):

yes

OpenStudy (abdullah1995):

so up till now is there any thing that you don't understand ?

OpenStudy (anonymous):

Okay then we do 7x dydx - 2y dydx = -4x-7y

OpenStudy (anonymous):

then we get -4x-7y/ (7x-2y)

OpenStudy (anonymous):

correcct?

OpenStudy (anonymous):

but the answer says the signs are suppose to be the opposite

OpenStudy (abdullah1995):

hmmm. that is really interesting. let me double check

OpenStudy (abdullah1995):

ohhh

OpenStudy (anonymous):

?

OpenStudy (abdullah1995):

our answer is right. your grading system multiplied everything my -1, to get rid of the -1 from the numerator ( top part of fraction)

OpenStudy (anonymous):

ohhhhhhhhhhh am I always suppose to do that| ?

OpenStudy (abdullah1995):

well, if you are giving a test on paper then you don't have to do that but since the grading system is built this way i guess you have to do it, yeah. ( i personally think it is pretty stupid)

OpenStudy (anonymous):

oh okay thanks ! so if its xy right, just put dydx beisde the x?

OpenStudy (abdullah1995):

in implicit differentiation you are differentiation y twice. the first time you differentiate it like normal. so for example if it xy^2 ==> (1)(y^2)+(x)(2y)(dy/dx)

OpenStudy (abdullah1995):

then after you differentiate it normally you add dy/dx on the y not the x

OpenStudy (abdullah1995):

so you are differentiating y twice in a way

OpenStudy (abdullah1995):

for practice try solving this one --> 2xy = xy^3 = 9

OpenStudy (anonymous):

okay so, 2x dydx + 2y = 3x^2 dydx + 3 y^2 = 0 ?

OpenStudy (abdullah1995):

let me see

OpenStudy (abdullah1995):

it was suppose to be a plus sign not equal sign. my mistake over there. the first = sign is a + sign

OpenStudy (anonymous):

oh okay, 2xy + xy^3 = 9, 2x dydx +y + 3x^2 dydx + 3y^2 = 0 ?

OpenStudy (abdullah1995):